OpenStudy (anonymous):

(x^15) divided (x^5) =

OpenStudy (anonymous):

So... :)

OpenStudy (anonymous):

if multiply causes exponents to add, what do you think dividing does?

OpenStudy (anonymous):

\[x^{15}/x^5=x^{15-5}\]

OpenStudy (anonymous):

x^10 ans....

OpenStudy (anonymous):

Easier example with x^4 / x^3: \[\frac{ x ^{4} }{ x ^{3} }=\frac{ x \times x \times x \times x }{ x \times x \times x } = x ^{1}\]